Alex has saved $252 at the bank. He wants to accumulate $1750 for a trip to soccer camp. What percent of his goal has been reached?
Answer by jim_thompson5910(35256)

Percentage = ( (Amount Saved)/(Total Needed) )*100 = (252/1750) * 100 = 0.144 * 100 = 14.4%


So he has reached 14.4% of his goal.